OpenStudy (anonymous):

Use the quadratic formula to solve. If needed , round to the nearest hundredth x^2 +3=-4x

pooja195 (pooja195):

Hi there :) Welcome to openstudy! Put the equation into standered form first \[\huge~\rm~\bf~ax^2+bx+c=0\]

pooja195 (pooja195):

How would we make x^2 +3=-4x equal to 0 ? Any ideas?

OpenStudy (anonymous):

let me try ... ummm \[x^2+3x+0+-4x?\]

OpenStudy (anonymous):

wow that awnser looks so dumb :( sorry..

pooja195 (pooja195):

Nope not quite....and we are all here to learn no worries :D x^2 +3=-4x Add 4x to both sides what do we get?

OpenStudy (anonymous):

\[4x+x^2+3=0\]

OpenStudy (anonymous):

right?

pooja195 (pooja195):

Almost! :) You have it just in the wrong order. \[\huge~\rm~\bf~x^2+4x+3=0 \] \[\huge~\rm~\bf~ax^2+bx+c=0\] What are the abc values? a=? b=? c=? Hint: Use the standered form to help you figure it out :)

OpenStudy (anonymous):

im drawling a blank right now

OpenStudy (anonymous):

do u mind if i add you onto a website that its a white board where you can show me?? sorry im more visual ..

OpenStudy (anonymous):

go here @pooja195 https://www.twiddla.com/2486992

pooja195 (pooja195):

\(\bbox[8pt, #99ff99,border:1px solid black]{\Large \Large{x=~} \huge{ \frac{-\color{magenta}{b} \pm\sqrt{ \color{magenta}{b} ^2-4 \color{blue}{a} \color{red}{c}}}{2 \color{blue}{a}} } ~ }\) when the equation is \(\Large\color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\). \(\Large\color{black}{ \color{blue}{1} x^2+ \color{magenta}{4}x+ \color{red}{3}=0 }\).

OpenStudy (anonymous):

okay so then ...you have to sub the 3?

OpenStudy (anonymous):

im sorry im homeschooled and my teacher doesnt do live lessons to show us how to do anything

pooja195 (pooja195):

I see. \(\bbox[8pt, #99ff99,border:1px solid black]{\Large \Large{x=~} \huge{ \frac{-\color{magenta}{4} \pm\sqrt{ \color{magenta}{4} ^2-4 \color{blue}{(1)} \color{red}{(3)}}}{2 \color{blue}{(1)}} } ~ }\) when the equation is \(\Large\color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\). \(\Large\color{black}{ \color{blue}{1} x^2+ \color{magenta}{4}x+ \color{red}{3}=0 }\).

pooja195 (pooja195):

^it's all color coordianted

pooja195 (pooja195):

Solve what's in the square root first \[\huge~\rm~\bf~4^2 -4(1)(3)\] DO NOT SQUARE ROOT THIS

OpenStudy (anonymous):

-4 x 1 = 4 -4 x 3 = -12

pooja195 (pooja195):

come onto the link you sent me :)

OpenStudy (anonymous):

ok